The average hospitality girls working in a certain area is 19 years old. A civic oriented group made a recent survey on the age of the hospitality girls working in the same locality. A random sample of 20 respondents showed an average of 15.4 years old with a standard deviation of 2.14 years old. Using a 0.05 level of significance, has the hospitality girls gone lower?

1
Expert's answer
2021-12-22T04:08:57-0500

Let "X" be a random variable representing the age of hospitality girls.

"n=20,\space \bar{x}=15.4,\space s=2.14"

The hypotheses tested are,

"H_0:\mu=19 \space vs\space H_1:\mu\lt 19"

The test statistic is given as,

"t={(\bar{x}-\mu)\over{s\over\sqrt{n}}}={15.4-19\over {2.14\over\sqrt{20}}}={-3.6\over0.4785}=-7.5232"

"t" is compared with the t distribution table value at "\alpha=0.05" with "n-1=20-1=19" degrees of freedom.

The table value is given as,

"t_{0.05,19}=-1.729133".

The null hypothesis is rejected if "t\lt t_{0.05,19}"

Since "t=-7.5232\lt t_{0.05,19}=-1.729133", we reject the null hypothesis and conclude that there is sufficient evidence to show that the age of hospitality girls has gone lower at 5% level of significance.